A triangle with an area of 54 has a base that is 3 times the height.
Write an equation for the area in terms of the height?
What is the length of the base?
What is the height?

Answer by Boreal(15235)

A=1/2 *b*h
but A=54 and b=3h
so 54=(1/2)3h^2
108=3h^2
h^2=36
h=6
base=18